OpenStudy (anonymous):

PLEASE HELP, IM SO CONFUSED. Factoring Polynomials 2x^2 + 15x - 38

OpenStudy (misssunshinexxoxo):

Simplifying 2x2 + 15x + -38 = 0 Reorder the terms: -38 + 15x + 2x2 = 0 Solve this ^

OpenStudy (anonymous):

15x + 2x^2=38?

OpenStudy (misssunshinexxoxo):

We had to reorder them Solving for variable 'x'. Factor a trinomial. (-19 + -2x)(2 + -1x) = 0

OpenStudy (anonymous):

how did you get that??

OpenStudy (misssunshinexxoxo):

Simplifying -19 + -2x = 0 Solving -19 + -2x = 0 -19 + 19 + -2x = 0 + 19 Combine like terms: -19 + 19 = 0 0 + -2x = 0 + 19 -2x = 0 + 19 Combine like terms: 0 + 19 = 19 -2x = 19 Divide each side by '-2'. x = -9.5 Simplifying x = -9.5

OpenStudy (anonymous):

thats not what the question is asking. I know the answer its (2x+19)(x-2) but idk how do get there

OpenStudy (misssunshinexxoxo):

2 + -1x = 0 Solving 2 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + -1x = 0 + -2 Combine like terms: 2 + -2 = 0 0 + -1x = 0 + -2 -1x = 0 + -2 Combine like terms: 0 + -2 = -2 -1x = -2 Divide each side by '-1'. x = 2 Simplifying x = 2